A chi square probability of 0.05 means what?
If your chi square test has a probability of 0.05 or less it is likely, but not certain, that your hypothesis is not correct.
What are two possible outcomes to prove an experiments hypothesis?
To prove the hypothesis.
To disprove the hypothesis.
The probability is 0.4231, approx.
How can a probability be estimated?
One way to estimate the probability of an event is to use a theoretical model to compare the relative likelihood of the event compared to all possible outcomes.
What does it mean for an event to have a probability of 1?
Probability is on a scale of 1 to 0 and 1 is that .an event will certainly happen whereas 0 is that an event will not happen
It is 1 - prob(it does not rain today and tomorrow)= 1 - prob(it does not rain today)*prob(it does not rain tomorrow)
= 1 - [1- prob(it does rain today)]*[1 - prob(it does rain tomorrow)]
= 1 - [1 - 0.8]*[1 - 0.5]
= 1 - 0.2*0.5
= 1 - 0.1 = 0.9
The probability is 0. One card cannot be a club and a spade!
What is the probability that the spinner will land 5?
The answer depends on the shape of the spinner and the numbers on it.
The answer will depend on what the disease is.
How do i find sample standard deviation from population standard deviation?
If the population standard deviation is sigma, then the estimate for the sample standard error for a sample of size n, is s = sigma*sqrt[n/(n-1)]
How many 7 card hands are possible with a 13 cards deck?
If the cards are all different then there are 13C7 = 1716 different hands.
It is 3/13.
When h has the value 5 what is 6h 9?
Unfortunately, limitations of the browser used by Answers.com means that we cannot see most symbols. It is therefore impossible to give a proper answer to your question. Please resubmit your question spelling out the symbols as "plus", "minus", "times", "divided by", "equals", "squared", "cubed" etc. There is no symbol between 6h and 9.
When you have 2 six sided dice what is the probability that the largest number rolled is a 3?
It is 1/6.
There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.
If you believe that the children's genders are not independent then you would need to get empirical evidence from all families with four or more children in which the first three children were girls. If there are g families in which the fourth is a girl and b where the fourth is a boy then the required probability is b/(g+b).
However, if you assume that the children's genders are independent events then, given that the probability of a boy is approx 0.52, the probability of the fourth child is a boy is 0.52
"Playing cards" are chosen at random.
Number of samples = 42C4 = 42*41*40*39/24 = 111930